### v3

@@ A.1 RSV-AZ-ST Model
×$\times\exp\left\{-\frac{\mu_{\lambda}\exp(-h_{t})}{2\lambda_{t}k(\delta)^{2}(1-\rho^{2})}\left[y_{t}-\sqrt{\frac{\lambda_{t}}{\mu_{\lambda}}}\frac{\exp(h_{t}/2)\delta(z_{0t}-c)}{\sqrt{1-c^{2}\delta^{2}}}\right.\right.$
@@
×$\times\exp\left\{-\frac{1}{2\sigma_{\eta}^{2}}(h_{t+1}-\mu-\phi(h_{t}-\mu))^{2}\right\},$
@@
×$\times\exp\left\{-\frac{1}{2}\sum_{t=1}^{T}h_{t}-\frac{\mu_{\lambda}\exp(-h_{T})}{2\lambda_{T}k(\delta)^{2}}\left[y_{T}-\sqrt{\frac{\lambda_{T}}{\mu_{\lambda}}}\frac{\exp(h_{T}/2)\delta(z_{0T}-c)}{\sqrt{1-c^{2}\delta^{2}}}\right]^{2}\right\}$
×$\times\exp\left\{-\sum_{t=1}^{T-1}\frac{\mu_{\lambda}\exp(-h_{t})}{2\lambda_{t}k(\delta)^{2}(1-\rho^{2})}\left[y_{t}-\sqrt{\frac{\lambda_{t}}{\mu_{\lambda}}}\frac{\exp(h_{t}/2)\delta(z_{0t}-c)}{\sqrt{1-c^{2}\delta^{2}}}\right.\right.$
@@
×$\times\exp\left\{-\frac{1}{2\sigma_{\eta}^{2}}\sum_{t=1}^{T-1}(h_{t+1}-\mu-\phi(h_{t}-\mu))^{2}-\frac{(1-\phi^{2})}{2\sigma_{\eta}^{2}}(h_{1}-\mu)^{2}\right\}$
×$\times\exp\left\{-\frac{1}{2}\sum_{t=1}^{T}z_{0t}^{2}I(z_{0t}>0)\right\}\times\prod_{t=1}^{T}\frac{\mu_{\lambda}^{1/2}(\nu/2)^{\nu/2}}{\Gamma(\nu/2)}\lambda_{t}^{-(\nu+1)/2-1}\exp\left(-\frac{\nu}{2\lambda_{t}}\right)$
×$\times\sigma_{u}^{-T}\times\exp\left\{-\frac{1}{2\sigma_{u}^{2}}\sum_{t=1}^{T}(x_{t}-\xi-h_{t})^{2}\right\}\times\pi(\bm{\theta}),$
@@ A.1.1 Generation of $\mu$
$\mu_{1}$ $=\sigma_{\mu 1}^{2}\left\{\sigma_{\mu 0}^{-2}\mu_{0}+\sigma_{\eta}^{-2}\left[(1-\phi^{2})h_{1}+\frac{1-\phi}{1-\rho^{2}}\sum_{t=1}^{T-1}\left(h_{t+1}-\phi h_{t}\right.\right.\right.$
@@ A.1.2 Generation of $\phi$
$\pi(\phi\mid\cdot)\propto$ $(1+\phi)^{a_{\phi 0}-1}(1-\phi)^{b_{\phi 0}-1}(1-\phi^{2})^{1/2}$
×$\times\exp\left\{-\sum_{t=1}^{T-1}\frac{\mu_{\lambda}\exp(-h_{t})}{2\lambda_{t}k(\delta)^{2}(1-\rho^{2})}\left[y_{t}-\sqrt{\frac{\lambda_{t}}{\mu_{\lambda}}}\cdot\frac{\exp(h_{t}/2)\delta(z_{0t}-c)}{\sqrt{1-c^{2}\delta^{2}}}\right.\right.$
@@
×$\times\exp\left\{-\frac{1}{2\sigma_{\eta}^{2}}\sum_{t=1}^{T-1}(h_{t+1}-\mu-\phi(h_{t}-\mu))^{2}-\frac{1-\phi^{2}}{2\sigma_{\eta}^{2}}(h_{1}-\mu)^{2}\right\}$
@@
×$\quad\times\exp\left\{-\sum_{t=1}^{T-1}\frac{\mu_{\lambda}\exp(-h_{t})}{2\lambda_{t}k(\delta)^{2}(1-\rho^{2})}\left[y_{t}-\sqrt{\frac{\lambda_{t}}{\mu_{\lambda}}}\cdot\frac{\exp(h_{t}/2)\delta(z_{0t}-c)}{\sqrt{1-c^{2}\delta^{2}}}\right.\right.$
@@
×$\quad\times\exp\left\{-\frac{1}{2\sigma_{\eta}^{2}}\sum_{t=1}^{T-1}(h_{t+1}-\mu-\phi(h_{t}-\mu))^{2}-\frac{1-\phi^{2}}{2\sigma_{\eta}^{2}}(h_{1}-\mu)^{2}\right\}.$
@@ A.1.3 Generation of $\delta$
$\pi(\delta\mid\cdot)\propto$ $(1+\delta)^{a_{\delta 0}-1}(1-\delta)^{b_{\delta 0}-1}\cdot k(\delta)^{-T}$
×$\times\exp\left\{-\frac{\mu_{\lambda}\exp(-h_{T})}{2\lambda_{T}k(\delta)^{2}}\left[y_{T}-\sqrt{\frac{\lambda_{T}}{\mu_{\lambda}}}\cdot\frac{\exp(h_{T}/2)\delta(z_{0T}-c)}{\sqrt{1-c^{2}\delta^{2}}}\right]^{2}\right\}$
×$\times\exp\left\{-\sum_{t=1}^{T-1}\frac{\mu_{\lambda}\exp(-h_{t})}{2\lambda_{t}k(\delta)^{2}(1-\rho^{2})}\left[y_{t}-\sqrt{\frac{\lambda_{t}}{\mu_{\lambda}}}\cdot\frac{\exp(h_{t}/2)\delta(z_{0t}-c)}{\sqrt{1-c^{2}\delta^{2}}}\right.\right.$
@@ A.1.6 Generation of $\nu$
$\pi(\nu\mid\cdot)\propto$ $\exp\left\{\frac{T}{2}\log\mu_{\lambda}+\frac{\nu T}{2}\log\left(\frac{\nu}{2}\right)-T\log\Gamma\left(\frac{\nu}{2}\right)-\frac{\nu}{2}\sum_{t=1}^{T}\left(\lambda_{t}^{-1}+\log\lambda_{t}\right)+(n_{\nu 0}-1)\log\nu-S_{\nu 0}\nu\right\}$
×$\times\exp\left\{-\sum_{t=1}^{T-1}\frac{\mu_{\lambda}\exp(-h_{t})}{2\lambda_{t}k(\delta)^{2}(1-\rho^{2})}\left[y_{t}-\sqrt{\frac{\lambda_{t}}{\mu_{\lambda}}}\cdot\frac{\exp(h_{t}/2)\delta(z_{0t}-c)}{\sqrt{1-c^{2}\delta^{2}}}\right.\right.$
@@
×$\times\exp\left\{-\frac{\mu_{\lambda}\exp(-h_{T})}{2\lambda_{T}k(\delta)^{2}}\left[y_{T}-\sqrt{\frac{\lambda_{T}}{\mu_{\lambda}}}\cdot\frac{\exp(h_{T}/2)\delta(z_{0T}-c)}{\sqrt{1-c^{2}\delta^{2}}}\right]^{2}\right\}.$
@@ A.1.8 Generation of $\bm{h}$
$\mu_{t}$ =$=\begin{cases}\left[\delta\bar{z}_{0t}+\sqrt{1-\delta^{2}}\dfrac{\rho}{\sigma_{\eta}}(\alpha_{t+1}-\phi\alpha_{t})\right]\sqrt{\lambda_{t}}\exp(\alpha_{t}/2)\gamma,&t=1,\ldots,n-1,\\ \delta\bar{z}_{0t}\sqrt{\lambda_{t}}\exp(\alpha_{t}/2)\gamma,&t=n,\end{cases}$ (71)
$\sigma_{t}^{2}$ =$=\begin{cases}(1-\rho^{2})(1-\delta^{2})\lambda_{t}\exp(\alpha_{t})\gamma^{2},&t=1,\ldots,n-1,\\ (1-\delta^{2})\lambda_{t}\exp(\alpha_{t})\gamma^{2},&t=n.\end{cases}$ (72)
@@ A.1.9 Generation of $\bm{\lambda}$
$\pi(\lambda_{t}|\cdot)\propto$ $\lambda_{t}^{-(\frac{\nu+1}{2}+1)}$
×$\times\exp\left\{-\frac{\nu}{2\lambda_{t}}-\frac{\mu_{\lambda}\exp(-h_{t})}{2\lambda_{t}k(\delta)^{2}(1-\rho^{2})}\left[y_{t}-\sqrt{\frac{\lambda_{t}}{\mu_{\lambda}}}\cdot\frac{\exp(h_{t}/2)\delta(z_{0t}-c)}{\sqrt{1-c^{2}\delta^{2}}}\right.\right.$
@@
$b_{t}$ =$=\begin{cases}\nu+\dfrac{\mu_{\lambda}\exp(-h_{t})y_{t}^{2}}{k(\delta)^{2}(1-\rho^{2})},&t=1,\ldots,n-1,\\[8.00003pt] \nu+\dfrac{\mu_{\lambda}\exp(-h_{t})y_{t}^{2}}{k(\delta)^{2}},&t=n,\end{cases}$
$g(\lambda_{t})$ =$=\begin{cases}\exp\left\{\sqrt{\dfrac{\mu_{\lambda}}{\lambda_{t}}}\cdot\dfrac{\exp(-h_{t}/2)y_{t}}{k(\delta)^{2}(1-\rho^{2})}\left[\dfrac{\delta(z_{0t}-c)}{\sqrt{1-c^{2}\delta^{2}}}+\rho\cdot\dfrac{k(\delta)}{\sigma_{\eta}}(h_{t+1}-\mu-\phi(h_{t}-\mu))\right]\right\},&t=1,\ldots,n-1,\\[10.00002pt] \exp\left\{\sqrt{\dfrac{\mu_{\lambda}}{\lambda_{t}}}\cdot\dfrac{\exp(-h_{t}/2)y_{t}}{k(\delta)^{2}}\cdot\dfrac{\delta(z_{0t}-c)}{\sqrt{1-c^{2}\delta^{2}}}\right\},&t=n.\end{cases}$
@@ A.2 RSV-FS-ST Model
×$\times\frac{1}{\sqrt{2\pi(1-\rho^{2})\sigma_{\eta}^{2}}}\exp\left\{-\frac{1}{2(1-\rho^{2})\sigma_{\eta}^{2}}\left[h_{t+1}-\mu-\phi(h_{t}-\mu)-\rho\sigma_{\eta}\exp(-h_{t}/2)y_{t}\right]^{2}\right\},$
@@
×$\times\exp\left\{-\frac{1}{2(1-\rho^{2})\sigma_{\eta}^{2}}\sum_{t=1}^{T-1}\left[h_{t+1}-\mu-\phi(h_{t}-\mu)-\rho\sigma_{\eta}\exp(-h_{t}/2)y_{t}\right]^{2}\right\}$
×$\times\exp\left\{-\frac{(1-\phi^{2})}{2\sigma_{\eta}^{2}}(h_{1}-\mu)^{2}\right\}\sigma_{u}^{-T}\exp\left\{-\frac{1}{2\sigma_{u}^{2}}\sum_{t=1}^{T}(x_{t}-\xi-h_{t})^{2}\right\}\pi(\bm{\theta}).$
@@ A.2.2 Generation of $\phi$
×$\times\exp\left\{-\frac{1}{2(1-\rho^{2})\sigma_{\eta}^{2}}\sum_{t=1}^{T-1}\left[h_{t+1}-\mu-\phi(h_{t}-\mu)-\rho\sigma_{\eta}\exp(-h_{t}/2)y_{t}\right]^{2}\right\}$
×$\times\exp\left\{-\frac{1-\phi^{2}}{2\sigma_{\eta}^{2}}(h_{1}-\mu)^{2}\right\}$
@@ A.2.3 Generation of $\rho$
×$\times\exp\left\{-\frac{1}{2(1-\rho^{2})\sigma_{\eta}^{2}}\sum_{t=1}^{T-1}\left[h_{t+1}-\mu-\phi(h_{t}-\mu)-\rho\sigma_{\eta}\exp(-h_{t}/2)y_{t}\right]^{2}\right\}.$
@@ A.2.4 Generation of $\sigma_{\eta}^{2}$
×$\times\exp\left\{-\frac{1}{2(1-\rho^{2})\sigma_{\eta}^{2}}\sum_{t=1}^{T-1}\left[h_{t+1}-\mu-\phi(h_{t}-\mu)-\rho\sigma_{\eta}\exp(-h_{t}/2)y_{t}\right]^{2}\right\}$
@@ A.2.9 Generation of $\bm{h}$
$g(h_{t}\mid\gamma,\nu)$ $=-\frac{\nu+1}{2}\log\left\{1+\frac{\left(\sigma_{*}y_{t}\exp(-h_{t}/2)+\mu_{*}\right)^{2}}{\nu}\left[\gamma^{-2}I\left(y_{t}\exp(-h_{t}/2)\geq-\frac{\mu_{*}}{\sigma_{*}}\right)\right.\right.$
@@
    ×$\times\exp\left\{-\frac{1}{2}h_{1}-\frac{1-\phi^{2}}{2\sigma_{\eta}^{2}}(h_{1}-\mu)^{2}-\frac{1}{2\sigma_{u}^{2}}(x_{1}-\xi-h_{1})^{2}\right\}.$
@@
    ×$\times\exp\left\{-\frac{1}{2(1-\rho^{2})\sigma_{\eta}^{2}}\left[h_{t}-\mu-\phi(h_{t-1}-\mu)-\rho\sigma_{\eta}\exp(-h_{t-1}/2)y_{t-1}\right]^{2}\right\}$
    ×$\times\exp\left\{-\frac{1}{2}h_{t}-\frac{1}{2\sigma_{u}^{2}}(x_{t}-\xi-h_{t})^{2}\right\}.$
@@
    ×$\times\exp\left\{-\frac{1}{2}h_{T}-\frac{1}{2\sigma_{u}^{2}}(x_{T}-\xi-h_{T})^{2}\right\}.$

### v4

@@ 5 Conclusion
An interesting avenue for future research is to extend the proposed framework to allow for alternative or multiple realized measures within the measurement equation. In the present study, the RSV specification relies on log-RV5 as a single realized proxy, while forecast evaluation is conducted using several realized measures, including RK, BV, and Med. Allowing the model to incorporate alternative realized measures, or a multi-measure RSV specification, could help assess the robustness of the reported gains with respect to the choice of realized volatility proxy and may also mitigate the distributional mismatch suggested by the non-normality of log-RV documented in the empirical analysis.
A related line of research is provided by score-driven volatility models with flexible innovation distributions, such as the GH skew-$t$-based SD–GARCH specification of Catania and Grassi (2022), as well as joint SD models for returns and realized volatility proposed by Catania and Proietti (2020). While these approaches require more elaborate dynamic updating rules and likelihood evaluations, they provide informative points of comparison and could offer further insights into the role of realized measures and distributional flexibility in volatility modeling.
@@ A.1 RSV-AZ-ST Model
[…]
@@
$\times\exp\left\{-\frac{1}{2\sigma_{\eta}^{2}}(h_{t+1}-\mu-\phi(h_{t}-\mu))^{2}\right\},$
@@
$\times\exp\left\{-\frac{1}{2}\sum_{t=1}^{T}h_{t}-\frac{\mu_{\lambda}\exp(-h_{T})}{2\lambda_{T}k(\delta)^{2}}\left[y_{T}-\sqrt{\frac{\lambda_{T}}{\mu_{\lambda}}}\frac{\exp(h_{T}/2)\delta(z_{0T}-c)}{\sqrt{1-c^{2}\delta^{2}}}\right]^{2}\right\}$
[…]
@@
$\times\exp\left\{-\frac{1}{2\sigma_{\eta}^{2}}\sum_{t=1}^{T-1}(h_{t+1}-\mu-\phi(h_{t}-\mu))^{2}-\frac{(1-\phi^{2})}{2\sigma_{\eta}^{2}}(h_{1}-\mu)^{2}\right\}$
$\times\exp\left\{-\frac{1}{2}\sum_{t=1}^{T}z_{0t}^{2}I(z_{0t}>0)\right\}\times\prod_{t=1}^{T}\frac{\mu_{\lambda}^{1/2}(\nu/2)^{\nu/2}}{\Gamma(\nu/2)}\lambda_{t}^{-(\nu+1)/2-1}\exp\left(-\frac{\nu}{2\lambda_{t}}\right)$
$\times\sigma_{u}^{-T}\times\exp\left\{-\frac{1}{2\sigma_{u}^{2}}\sum_{t=1}^{T}(x_{t}-\xi-h_{t})^{2}\right\}\times\pi(\bm{\theta}),$
@@ A.1.1 Generation of $\mu$
[…]
@@ A.1.2 Generation of $\phi$
$\pi(\phi\mid\cdot)\propto\$$(1+\phi)^{a_{\phi 0}-1}(1-\phi)^{b_{\phi 0}-1}(1-\phi^{2})^{1/2}$
[…]
@@
$\times\exp\left\{-\frac{1}{2\sigma_{\eta}^{2}}\sum_{t=1}^{T-1}(h_{t+1}-\mu-\phi(h_{t}-\mu))^{2}-\frac{1-\phi^{2}}{2\sigma_{\eta}^{2}}(h_{1}-\mu)^{2}\right\}$
@@
[…]
@@
$\quad\times\exp\left\{-\frac{1}{2\sigma_{\eta}^{2}}\sum_{t=1}^{T-1}(h_{t+1}-\mu-\phi(h_{t}-\mu))^{2}-\frac{1-\phi^{2}}{2\sigma_{\eta}^{2}}(h_{1}-\mu)^{2}\right\}.$
@@ A.1.3 Generation of $\delta$
$\pi(\delta\mid\cdot)\propto\$$(1+\delta)^{a_{\delta 0}-1}(1-\delta)^{b_{\delta 0}-1}\cdot k(\delta)^{-T}$
$\times\exp\left\{-\frac{\mu_{\lambda}\exp(-h_{T})}{2\lambda_{T}k(\delta)^{2}}\left[y_{T}-\sqrt{\frac{\lambda_{T}}{\mu_{\lambda}}}\cdot\frac{\exp(h_{T}/2)\delta(z_{0T}-c)}{\sqrt{1-c^{2}\delta^{2}}}\right]^{2}\right\}$
[…]
@@ A.1.6 Generation of $\nu$
$\pi(\nu\mid\cdot)\propto\$$\exp\left\{\frac{T}{2}\log\mu_{\lambda}+\frac{\nu T}{2}\log\left(\frac{\nu}{2}\right)-T\log\Gamma\left(\frac{\nu}{2}\right)-\frac{\nu}{2}\sum_{t=1}^{T}\left(\lambda_{t}^{-1}+\log\lambda_{t}\right)+(n_{\nu 0}-1)\log\nu-S_{\nu 0}\nu\right\}$
[…]
@@
$\times\exp\left\{-\frac{\mu_{\lambda}\exp(-h_{T})}{2\lambda_{T}k(\delta)^{2}}\left[y_{T}-\sqrt{\frac{\lambda_{T}}{\mu_{\lambda}}}\cdot\frac{\exp(h_{T}/2)\delta(z_{0T}-c)}{\sqrt{1-c^{2}\delta^{2}}}\right]^{2}\right\}.$
@@ A.1.8 Generation of $\bm{h}$
$\mu_{t}$ $=\begin{cases}\left[\delta\bar{z}_{0t}+\sqrt{1-\delta^{2}}\dfrac{\rho}{\sigma_{\eta}}(\alpha_{t+1}-\phi\alpha_{t})\right]\sqrt{\lambda_{t}}\exp(\alpha_{t}/2)\gamma,&t=1,\ldots,n-1,\\ \delta\bar{z}_{0t}\sqrt{\lambda_{t}}\exp(\alpha_{t}/2)\gamma,&t=n,\end{cases}$ (82)
$\sigma_{t}^{2}$ $=\begin{cases}(1-\rho^{2})(1-\delta^{2})\lambda_{t}\exp(\alpha_{t})\gamma^{2},&t=1,\ldots,n-1,\\ (1-\delta^{2})\lambda_{t}\exp(\alpha_{t})\gamma^{2},&t=n.\end{cases}$ (83)
@@ A.1.9 Generation of $\bm{\lambda}$
$\pi(\lambda_{t}|\cdot)\propto\$$\lambda_{t}^{-(\frac{\nu+1}{2}+1)}$
[…]
@@
$b_{t}$ $=\begin{cases}\nu+\dfrac{\mu_{\lambda}\exp(-h_{t})y_{t}^{2}}{k(\delta)^{2}(1-\rho^{2})},&t=1,\ldots,n-1,\\[8.00003pt] \nu+\dfrac{\mu_{\lambda}\exp(-h_{t})y_{t}^{2}}{k(\delta)^{2}},&t=n,\end{cases}$
$g(\lambda_{t})$ $=\begin{cases}\exp\left\{\sqrt{\dfrac{\mu_{\lambda}}{\lambda_{t}}}\cdot\dfrac{\exp(-h_{t}/2)y_{t}}{k(\delta)^{2}(1-\rho^{2})}\left[\dfrac{\delta(z_{0t}-c)}{\sqrt{1-c^{2}\delta^{2}}}+\rho\cdot\dfrac{k(\delta)}{\sigma_{\eta}}(h_{t+1}-\mu-\phi(h_{t}-\mu))\right]\right\},&t=1,\ldots,n-1,\\[10.00002pt] \exp\left\{\sqrt{\dfrac{\mu_{\lambda}}{\lambda_{t}}}\cdot\dfrac{\exp(-h_{t}/2)y_{t}}{k(\delta)^{2}}\cdot\dfrac{\delta(z_{0t}-c)}{\sqrt{1-c^{2}\delta^{2}}}\right\},&t=n.\end{cases}$
@@ A.2 RSV-FS-ST Model
$\times\frac{1}{\sqrt{2\pi(1-\rho^{2})\sigma_{\eta}^{2}}}\exp\left\{-\frac{1}{2(1-\rho^{2})\sigma_{\eta}^{2}}\left[h_{t+1}-\mu-\phi(h_{t}-\mu)-\rho\sigma_{\eta}\exp(-h_{t}/2)y_{t}\right]^{2}\right\},$
@@
$\times\exp\left\{-\frac{1}{2(1-\rho^{2})\sigma_{\eta}^{2}}\sum_{t=1}^{T-1}\left[h_{t+1}-\mu-\phi(h_{t}-\mu)-\rho\sigma_{\eta}\exp(-h_{t}/2)y_{t}\right]^{2}\right\}$
$\times\exp\left\{-\frac{(1-\phi^{2})}{2\sigma_{\eta}^{2}}(h_{1}-\mu)^{2}\right\}\sigma_{u}^{-T}\exp\left\{-\frac{1}{2\sigma_{u}^{2}}\sum_{t=1}^{T}(x_{t}-\xi-h_{t})^{2}\right\}\pi(\bm{\theta}).$
@@ A.2.2 Generation of $\phi$
$\times\exp\left\{-\frac{1}{2(1-\rho^{2})\sigma_{\eta}^{2}}\sum_{t=1}^{T-1}\left[h_{t+1}-\mu-\phi(h_{t}-\mu)-\rho\sigma_{\eta}\exp(-h_{t}/2)y_{t}\right]^{2}\right\}$
$\times\exp\left\{-\frac{1-\phi^{2}}{2\sigma_{\eta}^{2}}(h_{1}-\mu)^{2}\right\}$
@@ A.2.3 Generation of $\rho$
$\times\exp\left\{-\frac{1}{2(1-\rho^{2})\sigma_{\eta}^{2}}\sum_{t=1}^{T-1}\left[h_{t+1}-\mu-\phi(h_{t}-\mu)-\rho\sigma_{\eta}\exp(-h_{t}/2)y_{t}\right]^{2}\right\}.$
@@ A.2.4 Generation of $\sigma_{\eta}^{2}$
$\times\exp\left\{-\frac{1}{2(1-\rho^{2})\sigma_{\eta}^{2}}\sum_{t=1}^{T-1}\left[h_{t+1}-\mu-\phi(h_{t}-\mu)-\rho\sigma_{\eta}\exp(-h_{t}/2)y_{t}\right]^{2}\right\}$
@@ A.2.9 Generation of $\bm{h}$
[…]
@@
    $\times\exp\left\{-\frac{1}{2}h_{1}-\frac{1-\phi^{2}}{2\sigma_{\eta}^{2}}(h_{1}-\mu)^{2}-\frac{1}{2\sigma_{u}^{2}}(x_{1}-\xi-h_{1})^{2}\right\}.$
@@
    $\times\exp\left\{-\frac{1}{2(1-\rho^{2})\sigma_{\eta}^{2}}\left[h_{t}-\mu-\phi(h_{t-1}-\mu)-\rho\sigma_{\eta}\exp(-h_{t-1}/2)y_{t-1}\right]^{2}\right\}$
    $\times\exp\left\{-\frac{1}{2}h_{t}-\frac{1}{2\sigma_{u}^{2}}(x_{t}-\xi-h_{t})^{2}\right\}.$
@@
    $\times\exp\left\{-\frac{1}{2}h_{T}-\frac{1}{2\sigma_{u}^{2}}(x_{T}-\xi-h_{T})^{2}\right\}.$